# Missing Pieces: How Framing Uncertainty Impacts Longitudinal Trust in AI Decision Aids — A Gig Driver Case Study

### v1

@@ Abstract.
Decision aids based on artificial intelligence (AI) are becoming increasingly common. When such systems are deployed in environments with inherent uncertainty, following AI-recommended decisions may lead to a wide range of outcomes. In this work, we investigate how the framing of uncertainty in outcomes impacts users’ longitudinal trust in AI decision aids, which is crucial to ensuring that these systems achieve their intended purposes. More specifically, we use gig driving as a representative domain to address the question: how does exposing uncertainty at different levels of granularity affect the evolution of users’ trust and their willingness to rely on recommended decisions? We report on a longitudinal mixed-methods study where we measured the trust of gig drivers as they interacted with an AI-based schedule recommendation tool. Statistically significant quantitative results indicate that participants’ trust in and willingness to rely on the tool for planning depended on the perceived accuracy of the tool’s estimates; that providing ranged estimates improved trust; and that increasing prediction granularity and using hedging language improved willingness to rely on the tool even when trust was low. Additionally, we report on interviews with participants which revealed a diversity of experiences with the tool, suggesting that AI systems must build trust by going beyond general designs to calibrate the expectations of individual users.
@@ 1. Introduction
Typically, decision aids based on artificial intelligence (AI) function by both recommending decisions and estimating how good the outcomes of following those decisions will be. AI decision aids are finding increasing adoption in all aspects of society, not just for organisations in healthcare (Wang et al., 2021), finance (Milana and Ashta, 2021), education (McConvey et al., 2023), commerce (Riahi et al., 2021), criminal justice (Završnik, 2019), and other domains, but also for the communication, expression, and productivity of individuals (Skjuve et al., 2023). Yet, when users follow the decisions of such systems, they can encounter unexpected variability in outcomes, including unintentional harms (McGregor, 2021). This is less likely to be caused by malice than by the uncertainty that inherently results from decision aids’ complex interactions with other people and systems (Russell et al., 2017).
How, then, can we ensure that AI decision aids are able to function in concert with their users to achieve their goal of helping make better decisions? This goal can only be achieved by moving beyond system-centric objectives to also focus on the perceptions and attitudes of end-users (Bansal et al., 2019). In particular, trust has been consistently recognised as a factor that mediates the interactions between AI systems and their users (Lockey et al., 2021). We follow the common definition of trust in AI as “the willingness of a person to be vulnerable to the actions of an AI decision aid, based on the expectation that it will perform a decision-making task important to the trustor” (Mayer et al., 1995; Solberg et al., 2022). It is well-known that the design of AI systems impacts the expectations and thus the trust of users (Burgess et al., 2023; Danner et al., 2020; Jacobs et al., 2021; Kocielnik et al., 2019; Kunkel et al., 2019; Scharowski et al., 2023).
As defined above, trust in an AI decision aid is shaped by context. Specifically, psychological models predict that user have a baseline level of trust in an AI decision aid, the expression of which is modulated by their perception of uncertainty in the outcomes of relying upon the system with some context (Solberg et al., 2022). Thus, the presence of uncertainty is important for ensuring that empirical measurements of users’ trust in AI are ecologically valid. However, for practical or ethical reasons, most empirical evaluations of how trust is impacted by the design of AI decision aids have relied on hypothetical (e.g., medicine) or low-stakes (e.g., video recommendation) deployment contexts (Ueno et al., 2022) devoid of such uncertainty.
In this work, we focus on the deployment context of gig driving, where drivers use their personal vehicles to fulfill ridesourcing and food delivery requests from platforms such as Uber, Lyft, DoorDash, and Instacart. This context combines environmental uncertainty with repeated interactions, as drivers must repeatedly make decisions about which requests to fulfill based on an incomplete view of the platform’s supply and demand. Within this context, we contribute an in situ case study, where we evaluate the effect of different framings of uncertainty on user trust in a AI decision aid. To do so, we deployed a schedule recommendation tool that also provides estimates of potential earnings, and measured gig drivers’ trust in the tool over longitudinal trials. Our goal is to address the following research questions with our findings:
@@
How do users’ perceptions of an AI decision aid’s accuracy in estimating outcomes impact their trust in it?
@@
How do users’ perceptions of an AI decision aid’s accuracy in estimating outcomes interact with their trust to affect their reliance on it?
[…]
How does the framing of uncertain predictions impact users’ trust and reliance in an AI decision aid?
@@ 2.1. Trust in AI
Different definitions of trust in AI have been proposed (Gillath et al., 2021; Kaplan et al., 2023; Lee and See, 2004; Mayer et al., 1995; Mcknight et al., 2011). We adopt a general definition based on (Mayer et al., 1995), whose model of organisational trust has been widely adopted by work on human-AI trust (Cai et al., 2019a; Liao and Sundar, 2022b; Meske and Bunde, 2020; Saßmannshausen et al., 2021; Textor et al., 2022). Specifically, we draw from a refinement of their model by (Solberg et al., 2022) that adapts it to the context of AI decision aids. An alternative but also widely-adopted (Aoki, 2021; Nourani et al., 2021; Rader et al., 2018) definition is that of (Lee and See, 2004). Both (Lee and See, 2004; Mayer et al., 1995) distinguish trust (an internal attitude) from reliance (the external, behavioural expression of that attitude). However, (Lee and See, 2004)’s definition is less general, as it entails an expectation of positive outcomes from following an AI system. (Solberg et al., 2022) also distinguishes reliance (allowing the AI to recommend decisions) from compliance (following the AI’s recommended decisions).
2.1.1. Measurement of Trust
Because trust is inherently contextual, no single consistent set of measures has been adopted by the literature. The reviews of (Kohn et al., 2021; Ueno et al., 2022) found a great diversity of measures for trust in automation, but also that experimental design choices were usually not informed by nor validated against models of trust. We help rectify this issue by drawing from (Solberg et al., 2022)’s model of trust in AI decision aids for our study design.
2.1.2. Trust in Challenging Contexts
Past work in human-AI interaction has assessed trust under a variety of contexts (Ueno et al., 2022), with some of them entailing unique challenges that impact trust dynamics. We focus on two such types of challenges in this work.
One challenge that impacts trust in AI is the level of uncertainty, which has also been referred to as risk (Stuck et al., 2021; Solberg et al., 2022). Uncertainty encapsulates different sources of variability that make it difficult for users to reason about the potential outcomes of relying on an AI; specifically, (Stuck et al., 2021) separated uncertainty into situational (i.e., inherent in the decision-making context) and relational (i.e., introduced by the AI) components. Prior studies on relational uncertainty in human-AI interaction have focused on medical (Burgess et al., 2023; Cai et al., 2019a; Beede et al., 2020; Jacobs et al., 2021; Wang et al., 2021; Sivaraman et al., 2023; Yang et al., 2023) and financial (Schoeffer et al., 2022; Scharowski et al., 2023) contexts, and generally relied on simulated data. In parallel to this work, the AI explainability literature has increasingly focused on trust calibration, which aims to give users realistic expectations regarding when and why AI systems may or may not perform well (Bhatt et al., 2021; Tomsett et al., 2020; Zhang et al., 2020). Various experiments have compared the effects of explanation types on trust calibration (Kocielnik et al., 2019; Bansal et al., 2021; Wang and Yin, 2021; Cau et al., 2023; Naiseh et al., 2023).
Another challenge that can dynamically impact trust in AI systems is their repeated interactions with users, the effects of which have received limited consideration in past work on human-AI interaction. (Ueno et al., 2022) reported that most studies in human-AI interaction evaluated measures of trust only once, although some studies have attempted to control temporal variation through randomisation (Danner et al., 2020; Naiseh et al., 2021; Wintersberger et al., 2021). Several studies have accounted for the evolution of trust over time more explicitly by comparing multiple measurements of trust over weeks or months (Jiang et al., 2004; Mou and Cohen, 2017; Mou et al., 2017; Kunkel et al., 2019).
Generally, studies of trust under these challenging contexts have relied on hypothetical or low-stakes scenarios. Even studies of deployed systems (Beede et al., 2020; Wang et al., 2021) have been limited to observational studies that did not compare multiple designs. Within our uncertain, longitudinal, and higher-stakes context, users have varying levels of experience with how environmental uncertainty impacts their earnings, but all of them are freshly experiencing the impacts of the AI’s relational uncertainty. Under this context, we contribute a mixed-methods study that compares the effects of different designs for an AI decision aid on trust. Specifically, we investigate a design dimension not explored by prior studies: how different framings of relational uncertainty in these recommendations using hedging language (Kaufman and Kirsh, 2022) can alter the longitudinal dynamics of trust.
2.2. Recommendation Systems
Recommendation systems are mediators of the interfaces between people and the technologies they use. However, these systems do not always serve to reveal users’ unstated preferences, but may instead influence them over repeated interactions (Gretzel and Fesenmaier, 2006). The dynamics of these interactions are a double-edged sword (Carroll et al., 2022): they may either lead to undesirable concentration (Ferraro et al., 2020; Zhou et al., 2021) or incentivise further exploration of content (Liang and Willemsen, 2022). There has been limited work on how these longitudinal interactions can potentially be shaped by the design of recommendation systems. It has, however, been found in single-shot settings that the specific wording used to frame recommendations can influence their perception by users. (Sundar et al., 2008; Liao and Sundar, 2022a) found that framing recommendations in terms of similarity to other users improves perceived quality and trust (i.e., the bandwagon effect). (Komiak and Benbasat, 2006; Sun et al., 2023; Suwanaposee et al., 2023) recovered conflicting conclusions on whether personalised framings of recommendations improve perceived quality (i.e., the Barnum effect). Our investigation of hedging addresses the question of how such effects can interact with users’ experiences to impact trust in recommendations longitudinally.
@@ 2.3. Gig Work and Gig Driving
Platform-based gig work is among the most prevalent contexts today in which AI-based systems make managerial decisions that control individuals (Holmström and Hällgren, 2022). Gig work offers workers the flexibility and autonomy to choose when and where they want to work (Lee et al., 2015). However, a growing body of work has demonstrated that this autonomy is hampered by the opacity of platforms’ assignment, pricing, and evaluation mechanisms (Griesbach et al., 2019; Lee et al., 2015; Singh et al., 2023; Tan et al., 2021; Wood et al., 2019). We follow a line of work that has responded to this issue by envisioning novel, worker-oriented designs of platform infrastructures (Stein et al., 2023; Zhang et al., 2022).
In particular, we tackle the challenge of information asymmetry in pricing mechanisms for gig driving platforms (Maffie, 2023; Rosenblat and Stark, 2016), which dynamically vary compensation based on consumer demand (Lee et al., 2015) and gamification (Zhang et al., 2022). Platforms have given drivers little visibility into the demand patterns and pricing structures that dictate their earnings (Griesbach et al., 2019; Lee et al., 2015), causing difficulties in planning and inequities in revenue (Miao et al., 2023). The closest-related works to our own are those of (Battifarano and Qian, 2019; Khan et al., 2022; Zhang et al., 2023), who designed tools that provide transparency into earnings. We go beyond these works by deploying such a tool as part of an AI decision aid that also gives schedule recommendations.
3. Schedule Recommendation Tool
In this section, we develop an AI-based schedule recommendation tool for gig drivers. To our knowledge, such a schedule recommendation tool has not been proposed in prior AI literature. We use this tool as an example AI decision aid for our study on the relationship between the framing of uncertainty in outcomes and the longitudinal trust in AI decision aids.
@@ 3.1. Decision Aid Design
We focus on one aspect of activity planning for gig drivers: choosing when to work. Most drivers seek to earn as much as they can from driving. However, the importance of this goal varies; some drivers rely on gig driving as a primary or sole source of income, while others treat it as a source of supplemental income (Ma and Hanrahan, 2019; Wu et al., 2019). The amount of time that drivers invest also varies; part-time drivers balance gig driving with other responsibilities, while others view it as a full-time commitment (Berger et al., 2020; Lee et al., 2015; Ma and Hanrahan, 2019). Our schedule recommendation tool is designed to help gig drivers decide when to work given their constraints and preferences. The tool consists of two modules.
First, an estimation module prospectively predicts , the earnings that drivers can expect during a specific hour on a specific weekday . These estimates could be computed by a machine learning model or directly averaged from historical data.
Second, a scheduling module takes the estimated earnings and each driver’s constraints as inputs to produce an optimal set of working times that leads to the highest possible earnings. This module solves a constrained optimisation problem:
where denotes hour on weekday , denotes whether time slot is recommended, is the expected earning from driving during slot from the estimation module, is a binary indicator of whether the driver is available during slot , is the maximum number of hours they wish to work on weekday , and is the maximum number of hours they wish to work per week. Following past work on gig driving (Griesbach et al., 2019; Lee et al., 2015; Maffie, 2023; Zhang et al., 2023), we focus on daily and weekly constraints for driving hours and earnings.
Some drivers who value earnings to a greater extent set personal targets for their hourly or daily earnings (Zhang et al., 2022). In this case, we propose to solve the following alternative problem:
where is the minimum amount the driver wishes to earn on weekday , and is the minimum amount the driver wishes to earn during the entire week.
@@ 3.2. Interface Design
Next, since our proposed AI decision aid needs to take into account each driver’s constraints, we sought to design a front-end interface that would allow drivers both to enter these constraints and to view the optimised schedule of working times generated by the tool based on them. Our design consists of a two-page interface.
@@
First, a constraint page (Figure S1 in Appendix D) elicits the constraints needed for Equation 1 or Equation 2 by prompting users to enter their availability and goals. The page first asks users to select an optimisation objective: whether they would like set “a maximum number of hours to drive” ( in Equation 1) or “a minimum amount of money to make” ( in Equation 2). Then, it asks users to choose to set daily ( and ) or weekly ( and ) constraints. To maximise users’ perceived control over the system (see Section 4.1), we allowed them to choose these options freely as opposed to enforcing them within conditions. Lastly, the page elicits availability information for each hour of each day ( in Equations 1 and 2) in a tabular format. Unlike (Zhang et al., 2023), we account for the fact that users’ availability may change between days.
@@
Second, a schedule page (Figure 1) presents a tabular schedule generated by solving Equation 1 or Equation 2 to the user. The table shows the estimated earning of each time slot in its corresponding cell; the solution is shown by highlighting the recommended time slots (i.e., the ones that lead to the highest earnings subject to constraints). A summary of the user’s expected earnings for the week is shown above the table. In the interests of maximising perceived control, we allowed users to revisit the constraint page until they were satisfied with the generated schedule.
4. Uncertainty and Trust
4.1. Research Hypotheses
Provided that all parameters are specified, Equations 1 and 2 can be easily solved to generate a recommended schedule. However, this problem entails environmental uncertainty: each time slot’s earnings will dynamically vary with demand and supply. Inaccurate earnings estimates could substantially impact the quality of the solution (Cameron et al., 2022). Furthermore, information asymmetry prevents drivers and our AI decision aid from observing the ground truth distribution of gigs in real time. Thus, further uncertainty is introduced by the AI system having to rely on historical data or other proxies. How does this uncertainty impact trust and reliance?
Regarding Research Questions 1 and 2, (Solberg et al., 2022)’s model proposes that the outcomes of reliance on an AI decision aid can impact users in two ways. First, they may affect perceptions of the system’s fundamental trustworthiness, in terms of performance (i.e., competency in tasks), process (i.e., transparency of behaviour), and purpose (helpfulness for beneficiaries) (Lee and Moray, 1992). Second, they may affect perceptions of the uncertainty present in a given context (situational risk), which modulates the expression of trust through reliance. Even if the user trusts an AI decision aid in general, they may still feel that trusting it in particularly uncertain contexts could lead to adverse outcomes that fall short of the AI’s promises. If the tool is able to accurately estimate the outcomes of users’ reliance, it may both enhance performance-based trust and lead to higher reliance in uncertain contexts. This leads us to the following two hypotheses:
[…]
Users who perceived the outcomes as meeting or exceeding estimates will have a greater level of performance-based trust than those who perceived the outcomes as falling below estimates.
[…]
Users with the same level of trust prior to interaction will exhibit higher reliance if they have perceived the outcomes as meeting or exceeding estimates.
Meanwhile, regarding Research Question 3, (Solberg et al., 2022)’s model predicts that the design of an AI decision aid impacts its trustworthiness. Which design dimensions are most likely to affect trust? (Holmström and Hällgren, 2022) identified two axes of AI system designs that define different types of human-AI interaction: algorithmic management, the extent of the AI’s control on tracking, evaluation, and decision-making; and transparency, the extent to which these processes can be understood. Along the axis of algorithmic management, (Solberg et al., 2022) hypothesised that high levels of perceived control over otherwise untrustworthy systems may lead to higher levels of trust (Ho et al., 2013). Along the axis of transparency, (Solberg et al., 2022) considered explainability methods as enhancing process-based trust, in accordance with empirical results (Ferrario and Loi, 2022; Shin, 2021), and increased visibility into error rates as harming performance-based trust. We believe that this latter design dimension modulates how users’ perceptions of uncertainty impact their trust and reliance. This leads us to the following two hypotheses:
[…]
Designs which present estimates with greater granularity but not error visibility increase users’ trust and reliance.
[…]
Designs which show ranged estimates decrease users’ trust and reliance through increasing error visibility.
4.2. Tool Conditions
Screenshot of a box titled “Shift Recommendation Tool” on the schedule page, showing Condition 1. An estimated weekly earning of $1 412 is highlighted with box A. Inset A shows variants of how the weekly earning is presented (Condition 2, a table showing “An average of…” with $210 for Monday, $194 for Tuesday, etc.; Conditions 3 and 4, a table showing “On an average week: $1 412”, “On a bad week: $391”, and “On a good week: $3 032”). A recommended schedule shows average earnings for each weekday and each hour. Two cells reading “Average: $15” and “Average: $12” are highlighted with box B. Inset B shows a variant of how the hourly earnings are presented (Conditions 3 and 4, cells reading “Average: $15, Low: $4, High: $34” and “Average: $12, Low: $4, High: $28”). Before the weekly estimate of $1 412, box C encloses text reading “Based on historical data, it is estimated that you will earn a weekly average of”. Inset C shows a variant of this text (Condition 4: “In general, every driver may encounter different conditions while they are driving, which leads to differences in earnings from individual to individual and from one day to the next. On average, based on historical data, a driver following this schedule will earn”).
In service of Research Question 3, we varied the design of the schedule page in our tool between 4 different framings of uncertainty. In 2 of the 4 designs, we accomplished this by presenting earning estimates as ranges instead of as a point estimate. The range consisted of 3 quantities: a mean earning ( for hourly earnings; for weekly earnings), a pessimistic earning (based on the 10th percentile of earnings), and an optimistic earning (based on the 90th percentile of earnings). Figure 1 shows the 4 conditions:

### v2

# Missing Pieces: How Do Designs that Expose Uncertainty Longitudinally Impact Trust in AI Decision Aids? An In Situ Study of Gig Drivers

@@ Abstract.
Decision aids based on artificial intelligence (AI) induce a wide range of outcomes when they are deployed in uncertain environments. In this paper, we investigate how users’ trust in recommendations from an AI decision aid is impacted over time by designs that expose uncertainty in predicted outcomes. Unlike previous work, we focus on gig driving — a real-world, repeated decision-making context. We report on a longitudinal mixed-methods study () where we measured gig drivers’ trust as they interacted with an AI-based schedule recommendation tool. Our results show that participants’ trust in the tool was shaped by both their first impressions of its accuracy and their longitudinal interactions with it; and that task-aligned framings of uncertainty improved trust by allowing participants to incorporate uncertainty into their decision-making processes. Additionally, we observed that trust depended on their characteristics as drivers, underscoring the need for more in situ studies of AI decision aids.
@@ 1. Introduction
The fairness, accountability, and transparency of algorithms are inextricably intertwined with the extents to which people are willing to interact with them in different contexts. Accordingly, the FAccT community has increasingly focused on studying trust in algorithmic systems, especially those based on artificial intelligence (AI) (Ferrario and Loi, 2022; Jacovi et al., 2021; Kim et al., 2023; Knowles and Richards, 2021; Liao and Sundar, 2022). In this paper, we study how people trust and rely on AI decision aids. AI decision aids function by (1) recommending decisions and (2) predicting how good the outcomes of following those decisions will be. Over recent years, AI decision aids have been adopted at a steadily increasing rate in a number of application domains (Milana and Ashta, 2021; McConvey et al., 2023; Riahi et al., 2021; Skjuve et al., 2023; Wang et al., 2021; Završnik, 2019). However, because AI decision aids are used in the context of people and other sociotechnical systems, interactions with these entities can lead to considerable uncertainty in outcomes (Russell et al., 2017; Macrae, 2021). When this uncertainty impacts the predictability of AI decision aids, users’ trust in the decision aid can be eroded (Jacovi et al., 2021; Solberg et al., 2022).
Prior literature on trust in AI decision aids under uncertainty can be organised into two complementary lines of work. One line of work has studied specific factors that influence trust in AI decision aids, using laboratory experiments in simulated, single-shot, and low-stakes scenarios that require limited domain expertise (Bansal et al., 2021; Cai et al., 2019; Jacobs et al., 2021; Kocielnik et al., 2019; Zhang et al., 2020). However, context is an important factor for trust in AI (Kim et al., 2023; Solberg et al., 2022). Therefore, the contrived nature of these experiments limits their generalisability to the real-world use contexts of AI decision aids. Another line of work has studied trust in real-world AI decision aids (Beede et al., 2020; Kim et al., 2023; Wang et al., 2021), using observational, qualitative studies to assess how existing users interact with these decision aids. Despite their ecological validity, these studies do not quantitatively assess design factors and thus provide limited insight into how new, trustworthy AI decision aids can be designed.
In this paper, we provide a deeper exploration of trust in AI decision aids by combining the strengths of these two lines of work. We contribute the first in situ study of how exposing the uncertainty of an AI decision aid has longitudinal impacts on users’ trust and reliance on the decision aid. Using the domain of gig driving — in which drivers use their personal vehicles to fulfil ridesourcing and food delivery requests from platforms such as Uber, Lyft, DoorDash, and Instacart — as a testbed, we study trust in a real-world, medium-to-high-stakes decision-making scenario where users have existing expertise. Specifically, we comparatively evaluate different designs that expose the potential for misprediction in an AI decision aid. We address the following research questions:
@@
How do users’ trust and reliance on an AI decision aid depend longitudinally on their perception of its predictive accuracy?
@@
How do different designs that expose the inherent uncertainty in predictive performance impact users’ trust and reliance on an AI decision aid?
We addressed these questions by conducting a longitudinal user study where gig drivers used an AI-based schedule recommendation tool. By measuring the trust and reliance of participants over repeated interactions, we tested the effects of exposing uncertainty in the tool’s predictions through range-based earnings estimates and hedging text. Our quantitative and qualitative findings show that participants’ initial perceptions of the tool’s accuracy improved their trust in it over time. In addition, range-based uncertainty not only improved trust and reliance in single-shot settings, but also strengthened it over repeated interactions; meanwhile, hedging text had the opposite effect.
@@ 2.1. Trust in AI
A lack of uniformity exists in the human-AI interaction literature on how to define and evaluate trust in AI systems (Ueno et al., 2022). We follow Mayer et al. (1995) in operationalising the constructs of trust and reliance: in the context of AI decision aids, they hypothesise that trust is “the willingness of a person to be vulnerable to the actions of an [AI decision aid], based on the expectation that it will perform a [decision-making task] important to the trustor”, and that reliance is the external, behavioural expression of that internal attitude (Solberg et al., 2022). In accordance with this definition, one point of broad consensus in the literature is that the design of AI systems impacts expectations of their performance in decision-making, and thus the trust of their users (Burgess et al., 2023; Danner et al., 2020; Jacobs et al., 2021; Kocielnik et al., 2019; Kunkel et al., 2019; Scharowski et al., 2023; Yu et al., 2019).
However, the majority of this work has been based on controlled laboratory experiments. Compared to real-world use contexts of AI decision aids, such experiments have two main limitations: (1) They are single-shot, involving only a single session of AI use with no temporal separation between decision points (Ueno et al., 2022). Several studies have compared two trust measurements (Mou and Cohen, 2017; Mou et al., 2017; Kunkel et al., 2019), but these studies still provide a limited perspective on the evolution of trust dynamics. (2) They are low-stakes, involving contrived or hypothetical decision-making scenarios in which an element of risk and thus vulnerability is largely absent (Kim et al., 2023). Some work has studied AI decision aids for decision-making domains entailing high stakes in the real world, such as the medical (Burgess et al., 2023; Cai et al., 2019; Beede et al., 2020; Jacobs et al., 2021; Wang et al., 2021; Sivaraman et al., 2023; Wysocki et al., 2023; Yang et al., 2023) and financial (Prabhudesai et al., 2023; Schoeffer et al., 2022; Scharowski et al., 2023) contexts. However, for practical and ethical reasons, participants in these studies cannot receive feedback from the real world for their decisions. A minority of work has evaluated trust in AI decision aids within their real-world contexts (Beede et al., 2020; Kim et al., 2023; Wang et al., 2021; Widder et al., 2021), but these have been limited to observational studies that did not compare multiple designs.
Uncertainty encapsulates sources of variability that make it difficult for users to reason about the outcomes of relying on an AI, thus increasing the risk of this reliance (Stuck et al., 2021). Various experimental studies have tested the effects of designs that make users more aware of the presence and impact of uncertainty on AI (Bansal et al., 2021; Cau et al., 2023; Kocielnik et al., 2019; Naiseh et al., 2023; Wang and Yin, 2021; Yang et al., 2024). This work has parallels to trust calibration in the AI explainability literature: giving users realistic expectations regarding when and why AI systems may or may not perform well (Bhatt et al., 2021; Tomsett et al., 2020; Zhang et al., 2020). In addition to the lack of real-world context for these experiments, one thus far underexplored design is the use of lexical hedging: verbiage that expresses uncertainty. Kim et al. (2024) assessed the effects of lexical hedging in a large language model’s medical answers on trust but not on reliance. Zhang et al. (2022b) measured trust and reliance on an AI decision aid with lexical hedging for a contrived shape identification task. We perform a real-world evaluation of two designs for presenting uncertainty in an AI decision aid: presenting scalar ranges for estimates, and qualifying estimates with lexical hedging.
@@ 2.2. Gig Work and Gig Driving
Gig work offers workers the flexibility and autonomy to choose when and where they want to work (Lee et al., 2015). However, this autonomy is hampered by the opacity of platforms’ assignment, pricing, and evaluation mechanisms (Griesbach et al., 2019; Lee et al., 2015; Singh et al., 2023; Tan et al., 2021; Wood et al., 2019). In particular, gig driving platforms impose information asymmetry (Maffie, 2023; Rosenblat and Stark, 2016) by dynamically varying their compensation mechanisms (Lee et al., 2015; Zhang et al., 2022a). For drivers, this lack of visibility leads to difficulties in planning (Griesbach et al., 2019) and inequities in revenue (Miao et al., 2023). This makes gig driving an exemplary context under which the impact of uncertainty, repeated interactions, and moderate financial stakes on AI decision aids can be assessed.
Within the gig driving setting, the most relevant prior work to our paper consists of studies that have designed predictive and prescriptive tools for gig drivers (Battifarano and Qian, 2019; Khan et al., 2022; Zhang et al., 2022a, 2023, 2024). Among these, Battifarano and Qian (2019); Khan et al. (2022) focused on evaluating the predictive accuracy of their systems and did not include a user design component; Zhang et al. (2022a, 2023) were formative studies that designed tools in collaboration with drivers but did not deploy them to evaluate users’ trust and reliance; and Zhang et al. (2024) tested the effects of different AI decision aid designs that displayed uncertainty, albeit using simulated taxi trip data and focusing on the decisions — not attitudes — of participants.
3. AI-Based Schedule Recommendation for Gig Driving
In this section, we describe an AI-based schedule recommendation tool for gig drivers. Like other AI decision aids (Solberg et al., 2022), this tool (1) recommends a set of decisions (i.e. a schedule) to achieve a set objective (see Section 3.1), and (2) predicts the outcomes (i.e. estimated earnings) of following those decisions. We use this tool as an exemplary AI decision aid to study the longitudinal relationship between the framing of outcome uncertainty and trust. The design of this tool was conceptualised and refined through a formative study, which we describe in Appendix B.
@@ 3.1. Decision Aid Design
We focus on one aspect of planning for gig drivers: choosing when to work, subject to their constraints and preferences. There is considerable variation in drivers’ habits along this dimension (Berger et al., 2020; Lee et al., 2015; Ma and Hanrahan, 2019). Choosing where to work is another key aspect of planning (Zhang et al., 2023), but we limited our study of uncertainty in this multi-objective problem to a single dimension.
Our tool consists of two modules. First, an estimation module prospectively predicts , the earnings that drivers can expect during a specific hour on a specific weekday . These could be computed by a machine learning model or averaged from historical data. Second, for each driver, a scheduling module uses the estimated earnings and the driver’s constraints as inputs to produce an optimal set of working times. To do so, it solves a constrained optimisation problem to set variables to 1 or 0, denoting whether the driver is recommended to work during each time slot .
Some drivers wish to maximise their earnings while minimising their driving hours. For these drivers, we maximise the objective function: the sum of the estimated earnings for all recommended timeslots from the estimation module, i.e. .
To set the constraints, we disallow timeslots when the driver is not available, i.e. where is an indicator of whether the driver is available during timeslot . We also place an upper bound on the total hours of recommended timeslots per day by , and per week by , i.e. .
Some drivers who value earnings to a greater extent set personal minimum targets for their hourly or daily earnings instead of restricting their driving hours. For these drivers, we minimise the objective function: the total hours of recommended timeslots throughout the week, i.e. .
To set the constraints, we disallow timeslots when the driver is not available, i.e. . We also place a lower bound on the estimated earnings per day by , and per week by , i.e. .
@@ 3.2. Interface Design
Next, we also designed a front-end interface that would allow drivers to interact with the tool. The interface was implemented as a HTML/CSS/JavaScript website using Django 4.1 (Django, 2023) and a PostgreSQL database. To mitigate potential biases, we designed our tool to be visually generic and distinct from apps or websites associated with any gig platforms.
@@
First, a constraint page (Figure 9 in Appendix F) elicits the constraints needed for the scheduling module. It prompts users to select an optimisation objective: whether to maximise earnings or minimise hours on a daily or weekly basis (, , , and ). To maximise perceived control over the tool, we allowed users to choose these options freely rather than assigning them as conditions. The page also elicits hourly availability information (). Unlike Zhang et al. (2023), we account for the fact that users’ availability may change between days.
@@
Second, a schedule page (Section 3.3) presents a tabular schedule to the user. The optimal schedule is shown by highlighting the recommended time slots, i.e. the ones that lead to the highest earnings. Again, in the interests of maximising perceived control over the tool, we allowed users to revisit the constraint page until they were satisfied with the schedule.
3.3. Interface Conditions
The schedule page uses the outputs of the estimation module to predict how much a driver following the recommended schedule would make per hour or per week. However, uncertainty inherently exists in these predictions, as they are based on historical data and their realisation is contingent upon which gigs are offered to and accepted by drivers. To address Research Question 2, we varied the design of the schedule page between four conditions (Figure 1):